What is the surface area of a room with the measurements of 6ft 11in.x12ft 11in. x11ft 5in ?
Answer by mananth(16946)

6ft 11in.x12ft 11in. x11ft 5in ?
surface area of a cuboid is given by the formula
2*L*B +2*B*H+ 2*L*H
6ft 11 in = 6 11/12 feet= 83/12 feet
..
12ft 11 in = 12 11/12 feet = 155/12 feet
..
11ft. 5 in = 11 5/12 feet = 137/12 feet
..
2*83/12*155/12 +2*155/12*137/12 +2*137/12*83/12
..
2*83*155 / 144 + 2*155*137/144 +2*137*83/144
..
LCM = 144
..
2*83*155 + 2*155*137 +2*137*83/144
=631.54 ft^2